OpenStudy (happyfeet8040):

WILL GIVE METAL, PLEASE HELP!!! Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (anonymous):

The equation of a line is \[y = mx + b\] What do the variables y, x, m, and b represent?

OpenStudy (anonymous):

give me head and ill help

OpenStudy (happyfeet8040):

b is the y-intercept, where the line crosses the y-axis and the slope is the number m that is multiplied on the x.

OpenStudy (bradely):

y= mx+c y=-3x+c -5=-3(2)+c c=1 y=-3x+1 Option D is correct Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

Since the slope is -3, what is a way that we can solve for the y-intercept of this line, since we have one point on the line?

OpenStudy (anonymous):

12

OpenStudy (happyfeet8040):

not sure